# Generating Diverse and High-Quality Texts by Minimum Bayes Risk Decoding

### v1

@@ Appendix C Additional Results
Table 7: Evaluation of the quality and diversity using various decoding algorithms on WMT’19 De-En and Ru-En dataset. The size of the output $k$ is set to 4. Hyperparameters of the sampling algorithms are epsilon $\epsilon=0.02$, top-k $k=10$, and nucleus $p=0.9$. The best score is bolded and the second best score is underlined.
@@
Quality Diversity Decoder min BLEU $\uparrow$ mean BLEU $\uparrow$ max BLEU $\uparrow$ Pairwise-BLEU $\downarrow$ distinct-1 $\uparrow$ distinct-2 $\uparrow$ distinct-3 $\uparrow$ WMT’19 De-En ($k=8$) Epsilon 13.94 27.32 43.54 33.94 0.25 0.40 0.48 Top-k 12.19 25.46 42.23 28.39 0.28 0.46 0.54 Nucleus 10.49 23.90 40.17 25.37 0.33 0.51 0.58 Ancestral 5.34 17.28 34.05 13.23 0.47 0.67 0.72 Beam 23.82 34.02 47.28 63.21 0.17 0.23 0.25 DBS-0.5 19.15 32.43 46.96 52.07 0.19 0.27 0.31 DBS-1.0 15.77 29.76 46.69 42.89 0.23 0.34 0.38 DBS-2.0 12.28 26.88 44.73 36.31 0.27 0.40 0.44 DBS-5.0 8.89 23.60 41.72 31.37 0.32 0.46 0.50 MBR 28.43 34.77 41.58 70.36 0.16 0.21 0.24 DMBR-0.1 25.98 34.60 43.78 63.20 0.17 0.24 0.27 DMBR-0.3 19.56 32.26 46.43 46.23 0.21 0.32 0.38 DMBR-0.5 11.67 26.52 45.01 27.75 0.28 0.46 0.54 DMBR-1.0 9.30 21.03 38.56 18.62 0.33 0.56 0.64 DMBR-2.0 9.05 20.14 36.95 18.01 0.33 0.56 0.64 KMBR 15.56 29.88 47.62 39.36 0.24 0.36 0.43 WMT’19 De-En ($k=12$) Epsilon 12.61 27.54 46.32 34.16 0.19 0.33 0.41 Top-k 10.93 25.56 45.03 28.68 0.22 0.39 0.48 Nucleus 9.30 23.83 42.75 25.41 0.27 0.45 0.53 Ancestral 4.09 16.77 36.51 12.61 0.42 0.64 0.70 Beam 21.69 33.32 49.22 60.45 0.13 0.18 0.21 DBS-0.5 15.74 30.82 48.86 46.99 0.15 0.24 0.28 DBS-1.0 12.24 27.83 48.15 37.74 0.19 0.30 0.36 DBS-2.0 8.89 24.52 45.74 30.21 0.23 0.37 0.43 DBS-5.0 5.56 20.34 42.22 23.72 0.29 0.46 0.52 MBR 26.39 34.50 43.50 67.08 0.12 0.17 0.20 DMBR-0.1 24.40 34.36 45.33 62.08 0.12 0.18 0.22 DMBR-0.3 18.57 32.45 48.22 48.08 0.15 0.24 0.30 DMBR-0.5 10.66 27.03 48.03 29.86 0.21 0.37 0.46 DMBR-1.0 8.76 21.68 41.96 20.44 0.25 0.46 0.56 DMBR-2.0 8.64 20.79 40.61 19.74 0.25 0.46 0.56 KMBR 12.30 28.36 49.59 35.44 0.19 0.32 0.39
@@
Table 8: Evaluation of the quality and diversity using various decoding algorithms on WMT’19 De-En dataset. The size of the output $k$ is set to 8 and 12. Hyperparameters of the sampling algorithms are $\epsilon=0.02$, $k=10$, and $p=0.9$ for epsilon sampling, top-k sampling, and nucleus sampling, respectively. The best score is bolded and the second best score is underlined.
@@
Quality Diversity Decoder min BLEU $\uparrow$ mean BLEU $\uparrow$ max BLEU $\uparrow$ Pairwise-BLEU $\downarrow$ distinct-1 $\uparrow$ distinct-2 $\uparrow$ distinct-3 $\uparrow$ WMT’19 Ru-En ($k=8$) Epsilon 13.38 25.95 41.34 34.92 0.24 0.39 0.47 Top-k 11.79 24.07 39.09 29.57 0.27 0.44 0.52 Nucleus 9.70 22.39 37.71 25.57 0.32 0.50 0.58 Ancestral 3.82 14.69 30.21 11.05 0.48 0.70 0.75 Beam 21.83 31.21 43.11 64.87 0.16 0.22 0.25 DBS-0.5 17.60 29.68 43.08 52.23 0.19 0.27 0.31 DBS-1.0 14.51 27.62 43.15 43.67 0.22 0.32 0.37 DBS-2.0 11.97 25.06 41.28 37.18 0.25 0.38 0.43 DBS-5.0 8.56 22.12 38.48 32.54 0.30 0.44 0.49 MBR 25.22 31.87 38.90 69.76 0.15 0.21 0.24 DMBR-0.1 23.01 31.65 40.82 62.81 0.16 0.23 0.27 DMBR-0.3 18.25 29.80 42.84 47.04 0.20 0.31 0.37 DMBR-0.5 11.26 24.99 42.27 28.25 0.26 0.44 0.53 DMBR-1.0 8.89 20.23 36.68 19.38 0.31 0.53 0.63 DMBR-2.0 8.68 19.51 35.20 18.85 0.31 0.54 0.63 KMBR 14.67 27.94 44.57 40.54 0.22 0.35 0.41 WMT’19 Ru-En ($k=12$) Epsilon 12.05 25.97 43.70 34.95 0.18 0.31 0.40 Top-k 10.28 24.05 41.80 29.48 0.21 0.37 0.46 Nucleus 8.27 22.46 40.72 25.60 0.26 0.44 0.52 Ancestral 2.95 14.68 33.05 11.06 0.43 0.66 0.72 Beam 20.33 30.77 45.10 62.37 0.12 0.17 0.20 DBS-0.5 14.61 28.47 45.04 47.34 0.15 0.23 0.27 DBS-1.0 11.61 25.96 44.89 38.50 0.18 0.29 0.35 DBS-2.0 8.71 22.96 42.24 30.84 0.22 0.36 0.42 DBS-5.0 5.17 19.23 38.88 24.43 0.27 0.45 0.51 MBR 23.63 31.49 40.51 66.12 0.11 0.17 0.20 DMBR-0.1 21.84 31.36 42.00 61.51 0.12 0.18 0.22 DMBR-0.3 17.36 29.96 44.47 48.76 0.14 0.23 0.29 DMBR-0.5 10.35 25.23 44.37 30.36 0.20 0.35 0.45 DMBR-1.0 8.29 20.82 40.27 21.31 0.23 0.43 0.54 DMBR-2.0 8.14 20.10 38.77 20.60 0.23 0.44 0.54 KMBR 11.88 26.69 46.17 36.90 0.18 0.30 0.38
@@
Table 9: Evaluation of the quality and diversity using various decoding algorithms on WMT’19 Ru-En dataset. The size of the output $k$ is set to 8 and 12. Hyperparameters of the sampling algorithms are $\epsilon=0.02$, $k=10$, and $p=0.9$ for epsilon sampling, top-k sampling, and nucleus sampling, respectively. The best score is bolded and the second best score is underlined.

### v2

@@ E.2 Summary of the Results
Table 9: Evaluation of the quality and diversity using various decoding algorithms on WMT’19 De-En and Ru-En dataset. The size of the output $k$ is set to 4. Hyperparameters of the sampling algorithms are epsilon $\epsilon=0.02$, top-k $k_{\mathrm{top}}=10$, and nucleus $p=0.9$. The best score is bolded and the second best score is underlined.
@@
Quality Diversity Decoder min BLEU $\uparrow$ mean BLEU $\uparrow$ max BLEU $\uparrow$ Pairwise-BLEU $\downarrow$ distinct-1 $\uparrow$ distinct-2 $\uparrow$ distinct-3 $\uparrow$ WMT’19 De-En ($k=8$) Epsilon 13.94 27.32 43.54 33.94 0.25 0.40 0.48 Top-k 12.19 25.46 42.23 28.39 0.28 0.46 0.54 Nucleus 10.49 23.90 40.17 25.37 0.33 0.51 0.58 Ancestral 5.34 17.28 34.05 13.23 0.47 0.67 0.72 Beam 23.82 34.02 47.28 63.21 0.17 0.23 0.25 DBS-0.5 19.15 32.43 46.96 52.07 0.19 0.27 0.31 DBS-1.0 15.77 29.76 46.69 42.89 0.23 0.34 0.38 DBS-2.0 12.28 26.88 44.73 36.31 0.27 0.40 0.44 DBS-5.0 8.89 23.60 41.72 31.37 0.32 0.46 0.50 MBR 28.43 34.77 41.58 70.36 0.16 0.21 0.24 DMBR-0.1 25.98 34.60 43.78 63.20 0.17 0.24 0.27 DMBR-0.3 19.56 32.26 46.43 46.23 0.21 0.32 0.38 DMBR-0.5 11.67 26.52 45.01 27.75 0.28 0.46 0.54 DMBR-1.0 9.30 21.03 38.56 18.62 0.33 0.56 0.64 DMBR-2.0 9.05 20.14 36.95 18.01 0.33 0.56 0.64 KMBR 15.56 29.88 47.62 39.36 0.24 0.36 0.43 WMT’19 Ru-En ($k=8$) Epsilon 13.38 25.95 41.34 34.92 0.24 0.39 0.47 Top-k 11.79 24.07 39.09 29.57 0.27 0.44 0.52 Nucleus 9.70 22.39 37.71 25.57 0.32 0.50 0.58 Ancestral 3.82 14.69 30.21 11.05 0.48 0.70 0.75 Beam 21.83 31.21 43.11 64.87 0.16 0.22 0.25 DBS-0.5 17.60 29.68 43.08 52.23 0.19 0.27 0.31 DBS-1.0 14.51 27.62 43.15 43.67 0.22 0.32 0.37 DBS-2.0 11.97 25.06 41.28 37.18 0.25 0.38 0.43 DBS-5.0 8.56 22.12 38.48 32.54 0.30 0.44 0.49 MBR 25.22 31.87 38.90 69.76 0.15 0.21 0.24 DMBR-0.1 23.01 31.65 40.82 62.81 0.16 0.23 0.27 DMBR-0.3 18.25 29.80 42.84 47.04 0.20 0.31 0.37 DMBR-0.5 11.26 24.99 42.27 28.25 0.26 0.44 0.53 DMBR-1.0 8.89 20.23 36.68 19.38 0.31 0.53 0.63 DMBR-2.0 8.68 19.51 35.20 18.85 0.31 0.54 0.63 KMBR 14.67 27.94 44.57 40.54 0.22 0.35 0.41
@@
Table 10: Evaluation of the quality and diversity using various decoding algorithms on WMT’19 De-En and Ru-En dataset. The size of the output $k$ is set to 8. Hyperparameters of the sampling algorithms are $\epsilon=0.02$, $k_{\mathrm{top}}=10$, and $p=0.9$ for epsilon sampling, top-k sampling, and nucleus sampling, respectively. The best score is bolded and the second best score is underlined.
@@
Quality Diversity Decoder min BLEU $\uparrow$ mean BLEU $\uparrow$ max BLEU $\uparrow$ Pairwise-BLEU $\downarrow$ distinct-1 $\uparrow$ distinct-2 $\uparrow$ distinct-3 $\uparrow$ WMT’19 De-En ($k=12$) Epsilon 12.61 27.54 46.32 34.16 0.19 0.33 0.41 Top-k 10.93 25.56 45.03 28.68 0.22 0.39 0.48 Nucleus 9.30 23.83 42.75 25.41 0.27 0.45 0.53 Ancestral 4.09 16.77 36.51 12.61 0.42 0.64 0.70 Beam 21.69 33.32 49.22 60.45 0.13 0.18 0.21 DBS-0.5 15.74 30.82 48.86 46.99 0.15 0.24 0.28 DBS-1.0 12.24 27.83 48.15 37.74 0.19 0.30 0.36 DBS-2.0 8.89 24.52 45.74 30.21 0.23 0.37 0.43 DBS-5.0 5.56 20.34 42.22 23.72 0.29 0.46 0.52 MBR 26.39 34.50 43.50 67.08 0.12 0.17 0.20 DMBR-0.1 24.40 34.36 45.33 62.08 0.12 0.18 0.22 DMBR-0.3 18.57 32.45 48.22 48.08 0.15 0.24 0.30 DMBR-0.5 10.66 27.03 48.03 29.86 0.21 0.37 0.46 DMBR-1.0 8.76 21.68 41.96 20.44 0.25 0.46 0.56 DMBR-2.0 8.64 20.79 40.61 19.74 0.25 0.46 0.56 KMBR 12.30 28.36 49.59 35.44 0.19 0.32 0.39 WMT’19 Ru-En ($k=12$) Epsilon 12.05 25.97 43.70 34.95 0.18 0.31 0.40 Top-k 10.28 24.05 41.80 29.48 0.21 0.37 0.46 Nucleus 8.27 22.46 40.72 25.60 0.26 0.44 0.52 Ancestral 2.95 14.68 33.05 11.06 0.43 0.66 0.72 Beam 20.33 30.77 45.10 62.37 0.12 0.17 0.20 DBS-0.5 14.61 28.47 45.04 47.34 0.15 0.23 0.27 DBS-1.0 11.61 25.96 44.89 38.50 0.18 0.29 0.35 DBS-2.0 8.71 22.96 42.24 30.84 0.22 0.36 0.42 DBS-5.0 5.17 19.23 38.88 24.43 0.27 0.45 0.51 MBR 23.63 31.49 40.51 66.12 0.11 0.17 0.20 DMBR-0.1 21.84 31.36 42.00 61.51 0.12 0.18 0.22 DMBR-0.3 17.36 29.96 44.47 48.76 0.14 0.23 0.29 DMBR-0.5 10.35 25.23 44.37 30.36 0.20 0.35 0.45 DMBR-1.0 8.29 20.82 40.27 21.31 0.23 0.43 0.54 DMBR-2.0 8.14 20.10 38.77 20.60 0.23 0.44 0.54 KMBR 11.88 26.69 46.17 36.90 0.18 0.30 0.38
@@
Table 11: Evaluation of the quality and diversity using various decoding algorithms on WMT’19 De-En and Ru-En dataset. The size of the output $k$ is set to 12. Hyperparameters of the sampling algorithms are $\epsilon=0.02$, $k_{\mathrm{top}}=10$, and $p=0.9$ for epsilon sampling, top-k sampling, and nucleus sampling, respectively. The best score is bolded and the second best score is underlined.
@@
Appendix F Pretrained Models and Codes used in the Experiments
We list the pretrained models and codes we used in the experiments in Table 16.
Text Generation Models WMT’19 (Section 4.1) Ng et al. (2019) https://github.com/facebookresearch/fairseq/blob/main/examples/wmt19/README.md MS COCO (Section 4.2) Li et al. (2023) https://huggingface.co/Salesforce/blip2-flan-t5-xl-coco SQuADv2 (Section 4.3) Tunstall et al. (2023) https://huggingface.co/HuggingFaceH4/zephyr-7b-beta CommonGen (Section 4.4) Tunstall et al. (2023) https://huggingface.co/HuggingFaceH4/zephyr-7b-beta XSum (Section 4.5) Lewis et al. (2020) https://huggingface.co/facebook/bart-large-xsum Models for Evaluation WMT’19 (Section 4.1) sacreBLEU: Post (2018) https://github.com/mjpost/sacrebleu MS COCO (Section 4.2) Sentence BERT: Song et al. (2020) https://huggingface.co/sentence-transformers/all-mpnet-base-v2 SQuADv2 (Section 4.3) Sentence BERT: Song et al. (2020) https://huggingface.co/sentence-transformers/all-mpnet-base-v2 CommonGen (Section 4.4) Sentence BERT: Song et al. (2020) https://huggingface.co/sentence-transformers/all-mpnet-base-v2 CommonGen (Section 4.4) Porter stemmer: Porter (1980) https://www.nltk.org/_modules/nltk/stem/porter.html
Table 16: List of pretrained models we used in the experiments.